Ahmer walks 795m south and 55m west 670m north find the displacement and distance?
1
Expert's answer
2020-03-10T11:28:56-0400

The distance


"d=795+55+670=1520\:\rm m."

The displacement


"s=\sqrt{(795-670)^2+55^2}=137\:\rm m."